OpenStudy (anonymous):

August hosted two dinner parties for his friends. Twenty-five guests attended the first party, and twenty-eight guests attended the second party. What is the percentage increase of the number of guests from the first party to the second party? 12% 90% 10% 88%

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (rainbow_dashie):

Oh, this, I am sorry. I suck at this

OpenStudy (anonymous):

what about ols?

OpenStudy (rainbow_dashie):

ya i can

OpenStudy (anonymous):

im goingt to close this so it can be me and u

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (anonymous):

hold on trying to get it!

OpenStudy (anonymous):

Akira estimated that a tree was 40 years old. The true age of the tree was 47 years old. Which expression can be used to calculate the percent error in Akira’s estimate? A. \[\frac{ 47-40 }{ 40 }\times 100%\] B.\[\frac{ 47-40 }{ 47 }\times 100%\] C.\[\frac{ 40-47 }{ 40 }\times 100%\] D.\[\frac{ 47-40 }{ 47 }\times 100%\]

OpenStudy (anonymous):

the last two they have bars in the front and the end

OpenStudy (anonymous):

this is 9:9

OpenStudy (rainbow_dashie):

dude, what is this??

OpenStudy (anonymous):

Math lol i will find something else lol

OpenStudy (anonymous):

did you do 9:10?

OpenStudy (anonymous):

@Rainbow_dashie